OpenStudy (anonymous):

A net force of 100 newton's is applied to a wagon for 5 seconds. This causes the wagon to undergo a change in momentum of..

OpenStudy (anonymous):

\[Ft=mv-mu=100\times5=500Ns\]

TheSmartOne (thesmartone):

Change in momentum is impulse Impulse=force x time Impulse =100 x 5 impulse=500 Your final answer is \[500Ns= 500\frac{ kg \times m }{ s^2}s= 500\frac{ kg \times m }{ s }\]